OpenStudy (anonymous):

What three points would you graph for -3x + 2y = 2

OpenStudy (anonymous):

Do you know how to find the slope? y=mx+b?

OpenStudy (anonymous):

No :(

OpenStudy (anonymous):

Okay, start by solving for y.

OpenStudy (anonymous):

How do I do that?

OpenStudy (anonymous):

-3x+2y=2 So you want to get y by itself.

OpenStudy (anonymous):

Add 3x to both sides, so you get 2y=3x+2

OpenStudy (anonymous):

Then in order to get y completely alone, divide everything by 2... So you have y=(3/2)x+1

OpenStudy (anonymous):

Are you with me so far?

OpenStudy (anonymous):

Kind of. But I'm confused. I do online school and this was never explained in the lesson... I reviewed it many times.

OpenStudy (anonymous):

Ah, okay. I'm homeschooled to, so I understand. Alright I'll start from the beginning.

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

In order to graph something you have to find the slope. The standard form of slope is y=mx+b

OpenStudy (anonymous):

So when you have-3x+2y=2, you have to put in in the form y=mx+b.

OpenStudy (anonymous):

With me so far?

OpenStudy (anonymous):

yes!

OpenStudy (anonymous):

Okay. So you need to solve for y. How do you get it alone? You do the opposite of the signs in the problem. So because you have -3x, you add it to both sides. Understand?

OpenStudy (anonymous):

Yes I do

OpenStudy (anonymous):

Alright, so you're left with 2y=3x+2. And then you divide by 2. Which is how we get the slope: y=(3/2)x+1

OpenStudy (anonymous):

Do you have any questions about how to do that part?

OpenStudy (anonymous):

No, I understand that much. But then I don't understand how I would get my points from that

OpenStudy (anonymous):

Okay.

OpenStudy (anonymous):

When you're graphing, the parts of the equation y=mx+b that you need to pay attention to are m and b.

OpenStudy (anonymous):

M is the slope, and b is the y-intercept. Catch me?

OpenStudy (anonymous):

oh, okay

OpenStudy (anonymous):

So first you plot b, which is one. Draw a dot one spot up on the y axis.

OpenStudy (anonymous):

Then, since the slope is 3/2, you go up three, and to the right two. Where does that land? That is your point.

OpenStudy (anonymous):

So 2,4 is my first point? Since I started at 1

OpenStudy (anonymous):

Is it? =) Let me check for you.

OpenStudy (anonymous):

Yes, you're right. So you need three points, correct?

OpenStudy (anonymous):

Correct

OpenStudy (anonymous):

Would I just continue to rise 3 and run two? or am i completely off? :/

OpenStudy (anonymous):

No, you're right. =) Now, sometimes you run out of room on your graph going up. So you can also go down. Do you know how to do that with your slope of 3/2?

OpenStudy (anonymous):

No, I don't

OpenStudy (anonymous):

Wait but just to make sure my three plots would be (3,4) (4,7) (6,10)

OpenStudy (anonymous):

omg i meant 2,4

OpenStudy (anonymous):

(2,4), not three

OpenStudy (anonymous):

yes! I meant to say that

OpenStudy (anonymous):

Okay. You've got it then. So as far as going down, you just turn the slope negative =) To go the opposite way, use the opposite sign. It's really simple =)

OpenStudy (anonymous):

Oh my gosh, thank you so much! I've been struggling with this for two days. You're a life saver!

OpenStudy (anonymous):

Haha, I totally understand how hard it can be when you don't have a teacher to explain it to you =) I'm glad I could help. Let me know if you have any more questions.

OpenStudy (anonymous):

Would you be able to help me through one more problem please?

OpenStudy (anonymous):

Sure, whatcha got?

OpenStudy (anonymous):

-x + 2y = 2

OpenStudy (anonymous):

Okay, so try putting it in the standard form of slope: y=mx+b.

OpenStudy (anonymous):

Do you know what to do first?

OpenStudy (anonymous):

Do I add?

OpenStudy (anonymous):

Mhmmm, good. Add x to both sides. So you would get....

OpenStudy (anonymous):

Uhhh 1x + 3y = 2? Or do I add to the 2? :/

OpenStudy (anonymous):

... What? Lol you lost me. Isn't the problem -x+2y=2?

OpenStudy (anonymous):

yes!

OpenStudy (anonymous):

If you add x, you should get 2y=x+2

OpenStudy (anonymous):

I'm really horrible at math and I don't learn very fast.

OpenStudy (anonymous):

It's okay! Do you understand how I get 2y=x+2?

OpenStudy (anonymous):

Yes I do! So you only add to one number then?

OpenStudy (anonymous):

Do you understand combining like terms?

OpenStudy (anonymous):

Kind of... I was hospitalized my entire freshman year and I missed half of the year of algebra. So I don't know a lot of things I should.

OpenStudy (anonymous):

But I feel like combining like terms kind of explains itself..

OpenStudy (anonymous):

Yes. So because there aren't any other x terms in the problem, you don't have a change in numbers.

OpenStudy (anonymous):

Okay, that makes sense

OpenStudy (anonymous):

All you're doing is moving x from the left side to the right side.

OpenStudy (anonymous):

So we have 2y=x+2. Do you know what to do next?

OpenStudy (anonymous):

divide?

OpenStudy (anonymous):

Yes. So if you divide each term by 2, what do you get?

OpenStudy (anonymous):

I'm not sure what I would do with x since it is only a variable

OpenStudy (anonymous):

When you have x, you should always assume there's a 1 in front of it. 1x is the same as x. -1x is the same as -x. Does that make sense?

OpenStudy (anonymous):

Oh, okay. Yeah it does!

OpenStudy (anonymous):

Okay, so if you divide each term by 2, what do you get?

OpenStudy (anonymous):

y = 2x + 1

OpenStudy (anonymous):

Not quite.

OpenStudy (anonymous):